A firm faces the demand function

P = 190 - 0.6Q and the total cost function

C = 40 + 30Q + 0.4Q2


1
Expert's answer
2022-02-17T04:24:28-0500

a)  What output will maximize profit?


"MC=p""MC=C'(q)=30+0.8q"


Then


"30+0.8q=190-0.6q""q=114"

b) What output will maximize total revenue?


"TR=pq""TR=190q-0.6q^2""TR'(q)=190-1.2q""TR'(q)=0=>190-1.2q=0""q=158"

c) What will the output be if the firm makes a profit of 4,760?


"TR-TC=190q-0.6q^2-(40+30q+0.4q^2)""=160q-q^2-40"


"160q-q^2-40=4760""q^2-160q+4800=0, q\ge 0""(q-80)^2=1600""q_1=80+40=40, q_2=80-40=40"